A and B are supplementary.

Find A and B

A= 6x-1

B = 5x-17

Answer by Earlsdon(6294)

A and B are supplementary so A+B = 180 degrees.
6x-1+(5x-17 ) = 180
11x-18 = 180
11x = 198
x = 18 so...
A = 6x-1
A = 6(18)-1
A = 108-1
A = 107 degrees.
B = 5x-17
B = 5(18)-17
B = 90-17
B = 73 degrees.